OpenStudy (anonymous):

What is the exact value of the square root of 18 over 3 the square root of 36.? Simplify if possible.

OpenStudy (anonymous):

is it multiplied or are these two separate problems? \[\sqrt{18/3}*\sqrt{36} or \sqrt{18/3} and \sqrt{36}\]

OpenStudy (anonymous):

yes i already got that,

OpenStudy (anonymous):

no i'm trying to understand which one is the one your trying to find the answer to. is it two seperate questions or one?